please help me solve this equation: p+5/6=1/2
Answer by checkley71(8403)

P+5/6=1/2
P=1/2-5/6
P=3/6-5/6
P=(3-5)/6
P=-2/6
P=-1/3
PROOF
-1/3+5/6=1/2
-2/6+5/6=1/2
3/6=1/2
1/2=1/2